Two dice are rolled. What is the probability of getting a sum of at least 5?
Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3817)


Here are the ways to get a sum of 4 or smaller
1+1 = 2
1+2 = 3
1+3 = 4

2+1 = 3
2+2 = 4

3+1 = 4

There are 3 items in the first block, 2 in the next, and 1 in the last.
There are 3+2+1 = 6 ways to achieve a sum of 4 or smaller.

There are 6*6 = 36 ways to roll two dice.
That gives 36-6 = 30 ways to get a sum of 5 or more. You can use a dice chart to verify.

Divide the values:
30/36 = 5/6 is the probability of getting a sum of 5 or larger.

Answer by ikleyn(52784)

.

On rolling dice problems see the lesson
    - Rolling a pair of fair dice
in this site.  You will find there many similar solved problems.


Read it and learn the subject from there.  After that,  you will be able easy solve such problems
ON  YOUR  OWN,  which is your  MAJOR  GOAL  visiting this forum  (I believe).


Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic  "Solved problems on Probability".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.